# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

### v1

@@ III-A2 Snippets of Standalone Functions
For the scenario of completing function bodies based on function signatures and docstrings, we used the pre-training annealing data, opc-annealing-corpus  [^7] https://huggingface.co/datasets/OpenCoder-LLM/opc-annealing-corpus/, from the OpenCoder project to construct function completion samples. This dataset contains over ten million lines of code snippets synthesized by GPT-4. Specifically, we extracted function implementations from complete code files to serve as the code snippets to be completed, while the function signature, docstring, and other external code related to the target function collectively formed the context. For Java, Python, C++, Go, JavaScript, and TypeScript, we used data from the synthetic_code_snippet subset. As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead. Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324  [^8] https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions.
@@ III-A3 Filtering Based on Data Quality
$\text{PPL}(x)=\exp(-\frac{1}{N}\sum_{t=1}^{N}\log P(x_{t}|x_{<t}))$
In this work, we compute the perplexity of the code completion and code infilling data on the LLM to be trained. The distribution of the perplexity values generally approximates a log-normal distribution. For the post-training process, data with very low perplexity represents knowledge the LLM has already mastered and can output correctly without further training. Conversely, data with very high perplexity deviates significantly from the distribution of the LLM’s base data, and this subset tends to have a higher proportion of erroneous data. We remove a certain percentage of the data with the lowest and highest perplexity to improve the overall quality of the training set [37, 38]. For the infilling data, we ultimately remove data outside of two standard deviations (2 $\sigma$) of the fitted log-normal distribution. For the completion data, we discard the lowest and highest 5% of the data.
@@ III-B Data Construction
The core of our data construction lies in deeply simulating the real-world coding habits and contextual needs of developers. Our method encompasses three key aspects: (1) For constructing the completion target, we combine fine-grained node extraction based on Abstract Syntax Trees (ASTs) with heuristic strategies that mimic user behavior. (2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs). (3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities
and suppress context repetition.
@@ III-B1 Inner File Context
Today, AI-driven coding tools like GitHub Copilot  [^10] https://github.com/features/copilot and Cursor  [^11] https://www.cursor.com have been widely adopted by the developer community. Because developers have certain inherent habits in their thinking and coding processes – habits which are in turn influenced and reshaped by these AI tools – the locations where they expect to trigger code completion tend to follow specific patterns. Based on this core idea, our work strives to construct code completion data that aligns as closely as possible with developer habits. We have designed multiple strategies for this purpose, which can be broadly categorized into two main types: a strategy based on Abstract Syntax Tree (AST) nodes (60% of the data) and a heuristic strategy based on user habits (40%).
@@
In this paper, we use the Tree-sitter  [^12] https://tree-sitter.github.io/tree-sitter parsing tool to parse code files into abstract syntax trees. Based on a variety of node extraction strategies, we select target nodes as the code snippets to be completed and extract their corresponding prefix and suffix contexts. Compared to previous related work [39, 11, 12], we categorize syntactically or functionally similar nodes together, thereby creating 13 more fine-grained data synthesis strategies. These are based on: classes and interfaces; methods; formal and actual parameters; imported modules; expressions; assignment operations and variable declarations; comments; conditional statements; loop statements; return statements; other common statements; annotations or decorators; and core concurrent programming statements (only in Go).
@@
  • Code Completion Following a Comment: When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line. To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
  • Code Completion Within Parentheses: Paired parentheses often enclose relatively independent code snippets. In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between. We sample data to replicate this common scenario.
  • Completion of Incomplete Single Lines: When a developer pauses typing and there is no code to the right of the cursor, code completion is often triggered to finish the current line. We designed two strategies for this scenario.
    • Completion based on a random position: we randomly select a point that is not at the beginning or end of a line and designate the rest of the line as the completion target.
    • Completion based on special symbols: we compile a list of specific keywords and special characters for each language, search for these tokens during construction, and use the remainder of the line as the target.
@@ III-C1 Finetune Stage
Training Parameters and Cost
In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup. The training data selected was the top 30% (k=0.3) of the original dataset sorted by complexity, and the input sequences were uniformly truncated to within 8000 tokens. The entire process occupied 4 H800 80G PCIE GPUs, with a global batch size of 64 (micro-batch size of 1). Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency.
@@ III-C2 Alignment Stage
Training Parameters and Cost
In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64. The model was trained for one epoch with a maximum learning rate of 5e-5, a linear learning rate warmup of 48 steps, and a cosine learning rate decay scheduler. The maximum length for training data was 8192. For LoRA, the rank and alpha were set to 32 and 64, respectively, and the LoRA training modules included various modules within Self-Attention and the feed-forward networks. The training was conducted on 4 NVIDIA H800 80G PCIE GPUs using DeepSpeed ZeRO-2 and a CPU offloading strategy. The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 1. It can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
@@ IV Evaluation
We conducted a comprehensive and rigorous evaluation of SynthCoder and other baseline models on both file-level and repository-level code completion tasks. Due to constraints in computational resources and time, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025. Models released earlier, such as CodeLlama-7b-hf  [^13] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded. Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder  [^14] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B. All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1  [^15] https://github.com/vllm-project/vllm. The final set of models used for comparison is as follows.
  • Qwen2.5-Coder-7B  [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-7B  A 7B parameter code language model developed by Alibaba, designed for code understanding and generation, serving as the base model in its series.
  • Qwen2.5-Coder-7B-Instruct  [^17] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct  The instruction-tuned version of the Qwen2.5-Coder-7B model.
  • deepseek-coder-6.7b-base  [^18] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base A 6.7B parameter base code model developed by DeepSeek, serving as the base model in its series.
  • deepseek-coder-6.7b-instruct  [^19] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct The instruction-tuned version of deepseek-coder-6.7b-base.
  • Qwen3-Coder-30B-A3B-Instruct  [^20] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct An efficient code model from Alibaba’s Qwen3 series, which employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
  • Seed-Coder-8B-Base  [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base An 8B parameter base code model developed by the ByteDance Seed team. The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities.
  • Seed-Coder-8B-Instruct  [^22] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct The instruction-tuned version of Seed-Coder-8B-Base.
  • DeepSeek-Coder-V2-Lite-Base  [^23] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base A lightweight version of DeepSeek-Coder-V2  [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base/, serving as a base model. It also utilizes a Mixture-of-Experts (MoE) architecture.
  • DeepSeek-Coder-V2-Lite-Instruct  [^25] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct The instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
  • StarCoder2-7B  [^26] https://huggingface.co/bigcode/starcoder2-7b A 7B parameter code model developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages, with a focus on code generation tasks.
@@
  • aiXcoder-7b-v2  [^27] https://huggingface.co/aiXcoder/aixcoder-7b-v2 A 7B parameter code model designed to fully leverage long-context information, particularly excelling at cross-file code completion.
@@
  • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository  [^30] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context. Therefore, we only evaluated this model in the file-level code completion tasks;
@@ IV-A File Level Completion
We utilized the dataset  [^31] https://github.com/aixcoder-plugin/aiXcoder-7B/tree/main/aiXcoder_ex-tended_dataset provided by [11] to evaluate the model’s Fill-in-the-Middle (FIM) code completion performance on files from real-world open-source projects. The specific evaluation metrics used were Exact Match (EM) and Edit Similarity (ES). Our implementation is based on the reproduction project of Qwen Coder  [^32] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval. We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens (as tokenized by Qwen2.5-Coder-7B-Instruct) or where the generation length was greater than 64 tokens. During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
@@ IV-B3 ExecRepoBench
In our reproduction of the ExecRepoEval evaluation, we implemented a rule-based fix to correct indentation errors in the code generated by the models. We use Pass@1 and Edit Similarity (ES) to measure model performance. For each model, we employed its corresponding prompt format for completion, with standardized limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens. The temperature was set to 0.
@@ IV-C Errors of Wrongly Repeating Suffix or Prefix
Model SC* AX* ERE* CoLT
StarCoder2-7B 0.33% 1.38% 30.58% 1.87%
DS-Coder-6.7B-Inst 0.13% 1.62% 4.90% 1.98%
DS-Coder-V2-Lite-Inst 0.15% 0.64% 2.15% 2.19%
DS-Coder-6.7B-Base 0.13% 0.77% 4.73% 2.10%
DS-Coder-V2-Lite-Base 0.15% 1.00% 4.64% 2.11%
Qwen2.5-Coder-7B-Inst 0.13% 3.23% 1.03% 2.58%
Qwen3-Coder-Flash 0.10% 1.16% 0.26% 1.62%
Qwen2.5-Coder-7B 0.23% 0.69% 0.09% 1.51%
Seed-Coder-8B-Inst 0.31% $\uparrow$210% 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
Seed-Coder-8B-Base 0.17% $\uparrow$70% 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
SynthCoder-SFT 0.10% -0% 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
SynthCoder-DPO-wo-RS 0.11% $\uparrow$10% 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
SynthCoder-DPO-RS 0.10% 0.38% 0.09% 0.49%
    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
@@
Model SC* AX* ERE* CoLT
StarCoder2-7B 0.00% 0.49% 0.17% 0.04%
DS-Coder-6.7B-Inst 0.00% 0.43% 0.17% 0.03%
DS-Coder-V2-Lite-Inst 0.00% 0.41% 0.60% 0.08%
DS-Coder-6.7B-Base 0.00% 0.57% 0.34% 0.06%
DS-Coder-V2-Lite-Base 0.00% 0.52% 0.95% 0.08%
Qwen2.5-Coder-7B-Inst 0.00% 0.17% 0.43% 0.04%
Qwen3-Coder-Flash 0.00% 0.15% 0.09% 0.04%
Qwen2.5-Coder-7B 0.00% 0.18% 0.43% 0.04%
Seed-Coder-8B-Inst 0.00% -0% 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
Seed-Coder-8B-Base 0.00% -0% 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
SynthCoder-SFT 0.00% -0% 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
SynthCoder-DPO-wo-RS 0.00% -0% 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
SynthCoder-DPO-RS 0.00% 0.05% 0.09% 0.02%
    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively. The CrossCodeEval dataset is omitted from these tables because its ground-truth completions do not contain full lines, meaning the specific type of whole-line repetition we define does not occur. We observed that for models using either the Prefix-Suffix-Middle (PSM) FIM format (e.g., the Qwen-Coder and DeepSeek series) or the Suffix-Prefix-Middle (SPM) FIM format (e.g., the Seed-Coder series), the frequency of erroneously repeating the suffix is generally higher than that of repeating the prefix. The StarCoder2-7B model exhibits a particularly high repetition rate, reaching a suffix repetition rate of 30.58% on the ExecRepoEval benchmark. While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates. The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs. SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets. Interestingly, even without the inclusion of specialized repetition-suppression data, the context repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT. This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the dispreferred examples in the preference pairs for DPO training. After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior. It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.

### v2

@@ III-A2 Snippets of Standalone Functions
For the scenario of completing function bodies based on function signatures and docstrings, we used the pre-training annealing data, opc-annealing-corpus [^7] https://huggingface.co/datasets/OpenCoder-LLM/opc-annealing-corpus/, from the OpenCoder project to construct function completion samples. This dataset contains over ten million lines of code snippets synthesized by GPT-4. Specifically, we extracted function implementations from complete code files to serve as the code snippets to be completed, while the function signature, docstring, and other external code related to the target function collectively formed the context. For Java, Python, C++, Go, JavaScript, and TypeScript, we used data from the synthetic_code_snippet subset. As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead. Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324888https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions written in Go.
@@ III-A3 Filtering Based on Data Quality
$\text{PPL}(x)=\exp\Big{(}-\frac{1}{N}\sum_{t=1}^{N}\log P(x_{t}|x_{<t})\Big{)}$
[…]
Figure 1: Composition of the dataset by generation strategy.
@@ III-B Data Construction
The core of our data construction lies in deeply simulating the real-world coding habits and contextual needs of developers. Our method encompasses three key aspects: (1) For constructing the completion target, we combine fine-grained node extraction based on Abstract Syntax Trees (ASTs) with heuristic strategies that mimic user behavior. (2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs). (3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities and suppress context repetition.
@@ III-B1 Inner File Context
The widespread adoption of AI-driven coding tools has profoundly influenced the developer community. These tools interact with and shape developers’ inherent coding habits, leading to predictable patterns in the locations where they trigger code completion. Motivated by this observation, our work aims to construct a code completion dataset that closely mimics these real-world developer behaviors. To this end, we have designed multiple data generation strategies, which, as illustrated in Figure 1, are grouped into three primary categories: Strategies Based on AST Nodes (blue, abbreviated as AST), which constitute the majority at $66.89\%$; Strategies Based on User Behavior Patterns (orange, abbreviated as User), accounting for $22.56\%$; and Strategies Based on Random Single or Multiple Lines (green, abbreviated as Rand), which make up the remaining.
@@
In this paper, we use the Tree-sitter [^10] https://tree-sitter.github.io/tree-sitter parsing tool to parse code files into abstract syntax trees. Based on a variety of node extraction strategies, we select target nodes as the code snippets to be completed and extract their corresponding prefix and suffix contexts. To create fine-grained data, we identified 13 distinct strategies based on syntactically or functionally similar AST nodes, as detailed in the blue sections of Figure 1. These strategies vary in prevalence. The most significant contributions come from targeting locations based on Expressions ($10.66\%$), Other Common Statement Blocks ($9.90\%$), and Methods ($8.90\%$). Other key strategies address fundamental constructs such as assignment operations and variable declarations ($6.99\%$), parameters and arguments ($6.53\%$), and conditional statements ($5.43\%$). The remaining strategies cover more specific contexts, including annotations or decorators ($0.46\%$) and concurrent statements in Go ($0.18\%$), ensuring comprehensive coverage of syntactic structures.
@@
  • Incomplete Single Lines (14.78%): When a developer pauses typing and there is no code to the right of the cursor, code completion is often triggered to finish the current line. We designed two strategies for this scenario.
    • Random Intra-Line Completion: The completion is triggered at a randomly selected position within a line of code. The subsequent text on that line serves as the completion target.
    • Syntax-Aware Triggered Completion: The completion is triggered immediately following a language-specific, syntactically significant token. We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access). The remainder of the line after the token is designated as the completion target.
  • Code Fragments Within Parentheses (4.86%): Paired parentheses often enclose relatively independent code snippets. In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between. We sample data to replicate this common scenario.
  • Code Fragments After Comments (2.92%): When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line. To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
Strategies Based on Random Single or Multiple Lines
@@ III-C1 Finetune Stage
Training Details
In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup. The training data selected was the top 30% (k=0.3) of the original dataset sorted by complexity, and the input sequences were uniformly truncated to within 8000 tokens. The entire process occupied 4 H800 80G PCIE GPUs, with a global batch size of 64 (micro-batch size of 1). Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency. Finally, the process accounted for 800 steps and finished in 20 hours.
@@ III-C2 Alignment Stage
Training Details
In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64. The model was trained for one epoch with a maximum learning rate of 5e-5, a linear learning rate warmup of 48 steps, and a cosine learning rate decay scheduler. The maximum length for training data was 8192. For LoRA, the rank and alpha were set to 32 and 64, respectively, and the LoRA training modules included various modules within Self-Attention and the feed-forward networks. The training was conducted on 4 NVIDIA H800 80G PCIE GPUs using DeepSpeed ZeRO-2 and a CPU offloading strategy. The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 2. The alignment process lasted for 400 steps within 24 hours, and it can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
@@ IV Evaluation
We conducted a comprehensive and rigorous evaluation of SynthCoder and baseline models on both file-level and repository-level code completion tasks. Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025. Models released earlier, such as CodeLlama-7b-hf [^11] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded. Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder [^12] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B. All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1131313https://github.com/vllm-project/vllm. The final set of models used for comparison is as follows.
  • aiXcoder-7b-v2141414https://huggingface.co/aiXcoder/aixcoder-7b-v2
@@
    A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
  • StarCoder2-7B [^15] https://huggingface.co/bigcode/starcoder2-7b
    A 7B parameter code LLM developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages.
  • DeepSeek-Coder-6.7B (-base [^16] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^17] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct)
    A pair of 6.7B code LLMs developed by DeepSeek. The former is the base model, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
  • Qwen2.5-Coder-7B [^18] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^19] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct)
    A pair of 7B code LLMs developed by Alibaba, designed for code understanding and generation. Qwen2.5-Coder-7B is the base model, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version of Qwen2.5-Coder-7B.
  • Seed-Coder-8B (-Base [^20] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct)
    A pair of 8B parameter code LLMs developed by the ByteDance. The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
  • DeepSeek-Coder-V2-Lite (-Base [^22] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base, -Instruct [^23] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct)
    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base. Both of them utilize a Mixture-of-Experts (MoE) architecture, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
  • Qwen3-Coder-30B-A3B-Instruct [^25] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct
    An efficient instruction-tuned code model from Alibaba’s Qwen3 series, which also employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
@@
  • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository [^28] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context. Therefore, we only evaluated this model in the file-level code completion tasks.
@@ IV-A File Level Completion
We utilized the dataset [^29] https://github.com/aixcoder-plugin/aiXcoder-7B/tree/main/aiXcoder_ex-tended_dataset provided by [11] to evaluate the model’s Fill-in-the-Middle (FIM) code completion performance on files from real-world open-source projects. The specific evaluation metrics used were Exact Match (EM) and Edit Similarity (ES). Our implementation is based on the reproduction project of Qwen Coder [^30] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval. We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens or where the generation length was greater than 64 tokens. During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
@@ IV-B3 ExecRepoBench
In our reproduction of the ExecRepoEval evaluation, we implemented a rule-based fix to correct indentation errors in the code generated by the models. We use Pass@1 and Edit Similarity (ES) to measure model performance. For each model, we employed its corresponding prompt format for completion, with limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens. The temperature was set to 0.
@@ IV-C Errors of Wrongly Repeating Suffix or Prefix
Model AX* ERE* CoLT
StarCoder2-7B 1.38% 30.58% 1.87%
DS-Coder-6.7B-Inst 1.62% 4.90% 1.98%
DS-Coder-V2-Lite-Inst 0.64% 2.15% 2.19%
DS-Coder-6.7B-Base 0.77% 4.73% 2.10%
DS-Coder-V2-Lite-Base 1.00% 4.64% 2.11%
Qwen2.5-Coder-7B-Inst 3.23% 1.03% 2.58%
Qwen3-Coder-Flash 1.16% 0.26% 1.62%
Qwen2.5-Coder-7B 0.69% 0.09% 1.51%
Seed-Coder-8B-Inst 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
Seed-Coder-8B-Base 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
SynthCoder-SFT 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
SynthCoder-DPO-wo-RS 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
SynthCoder-DPO-RS 0.38% 0.09% 0.49%
    AX: aiXcoder; ERE: ExecRepoEval.
@@
Model AX* ERE* CoLT
StarCoder2-7B 0.49% 0.17% 0.04%
DS-Coder-6.7B-Inst 0.43% 0.17% 0.03%
DS-Coder-V2-Lite-Inst 0.41% 0.60% 0.08%
DS-Coder-6.7B-Base 0.57% 0.34% 0.06%
DS-Coder-V2-Lite-Base 0.52% 0.95% 0.08%
Qwen2.5-Coder-7B-Inst 0.17% 0.43% 0.04%
Qwen3-Coder-Flash 0.15% 0.09% 0.04%
Qwen2.5-Coder-7B 0.18% 0.43% 0.04%
Seed-Coder-8B-Inst 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
Seed-Coder-8B-Base 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
SynthCoder-SFT 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
SynthCoder-DPO-wo-RS 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
SynthCoder-DPO-RS 0.05% 0.09% 0.02%
    AX: aiXcoder; ERE: ExecRepoEval.
We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively. The CrossCodeEval dataset is omitted from these tables because its ground-truth completions do not contain full lines, meaning the specific type of whole-line repetition we define does not occur. We observed that for models using either the Prefix-Suffix-Middle (PSM) FIM format (e.g., the Qwen-Coder and DeepSeek series) or the Suffix-Prefix-Middle (SPM) FIM format (e.g., the Seed-Coder series), the frequency of erroneously repeating the suffix is generally higher than that of repeating the prefix. The StarCoder2-7B model exhibits a particularly high repetition rate, reaching a suffix repetition rate of 30.58% on the ExecRepoEval benchmark. While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates. Noteworthly, The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs. SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets. Interestingly, even without the inclusion of specialized repetition-suppression data, the repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT. This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the negative examples in the preference pairs for DPO training. After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior. It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.